OpenStudy (elisuzsmith):

Find all the zeros of the equation x^3 – x^2 + 5x – 5 = 0

OpenStudy (campbell_st):

well you can group in pairs \[x^3 - x^2 + 5x - 5 = (x^3 - x^2) + (5x - 5) = 0\] look at removing common factors from each pair and you may notice you will get a common factor which is a binomial... you should get a quadratic and linear binomials... hope this helps

OpenStudy (elisuzsmith):

thanks